The weights of 1,000 children average is 50 kg and the standard deviation is 5 kg. How many children weigh between 40 kg and 55 kg?


1
Expert's answer
2021-03-26T05:20:36-0400

Here the everage weights of the "1000" children is "50" kg with standard deviation "5" kg.

Let "X" be a random variable denotes the weight of the students.

Then "X" is normally distributed with mean "50" kg and standard deviation "5" kg.

Therefore we have "\mu =50" and "\sigma=5" .

Let us take "Z= \frac{X-\mu}{\sigma}" . Then "Z=\frac{X-50}{5}".

Now we have to find "P(40<X<55)."

"\therefore P(40<X<55)=P(\frac{40-50}{5}<Z<\frac{55-50}{5})"

"=P(-2<Z<1)"

"=P(0< Z< 2)+P(0< Z< 1)"

"=0.4772+0.3413" [ from normal distribution table ]

"=0.8185"

So the number of children weight between "40" kg and "55" kg is "=(1000\u00d70.8185)=819" (approximately)